Find the equation for the following Op-Amp

Twith R Lum & min SO
Answer #1

o ] given Rr I ISR. For ideal op-amp - V+ =V_ aus It= 1 = 0 Amp • For © . V+,= v* = V aus u. = V+ = Vy Now From circuit For -as IPy = IP Now U₂= V4 + P4 IR4 V = va fi + ]- v3 (91) Voz va protein) va (y SI 3. aus - U=V4 - Pq IP4 V = Vy - PM (Montag anlu,= V4 [P3-Py] + V3P4 R3 Now X Now For op-aup ② - U+;= Vzx Phot aut= q (Pst Py) - By V3 x R2 P3 RAP2 1. R2 Val 1 X 3 1 og beB-tyt P4-5R104 (P3TR4) - V3 Ry] Jl R3( Ri+P2) R[V4 (P3TR4) _V3 Ry] • -- BIR&R) Vi - R2 - Vo = Be TVA (P3T P4)-V3 Rii] Tuy (@g